a) Suppose that the depreciation rate of capital is 15% and the real interest rate is 5%. What is the desired level of the capital stock if the marginal product of capital is given by:

MPK = 1 - 0.16 K
Answer #1

The desired level of capital is the capital at which the firm maximizes its profit. So, a firm would make an investment when the marginal product of capital is equal to user cost of capital.

User cost of capital = Depreciation + Real rate of interest

MPK = User cost of capital

1 - 0.16*K = 0.15 +0.05

1-0.2 = 0.16*K

0.8 / 0.16 = K

K = 5

So, desired level of capital is 5.